OpenStudy (anonymous):

A car travels 100 meters in 5 seconds. What is the speed of the car? speed = distance over time; 1 mile = 1609 meters 8.9 miles per hour 16 miles per hour 20 miles per hour 44.7 miles per hour

OpenStudy (anonymous):

v = s / t = 100 m / 5 s = ? m/s 1 h = 3600 s 1 mile = 1609 m

OpenStudy (anonymous):

and this is chemistry department so you probably want to post this question in physics department